# Uncovering thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids

### v1

@@ Appendix A Holographic setup
To investigate holographic superfluids with spontaneously broken symmetry, we turn off electromagnetic fields and scalar sources at the boundary, meaning $a_{x}=a_{y}=(\Psi_{I})_{0}=0$. Consequently, the charge conservation equation at the boundary is given by $-\partial_{t}b_{t}=b_{i}$, where $-b_{t}=n$ represents the charge density and $b_{i}=j_{i}$ is the charge current. The superfluid velocity is determined by $\bm{v}^{s}=\nabla\theta|_{z=0}$, where $\theta$ is the phase (the Goldstone field) of the order parameter $\langle\mathcal{\hat{O}}\rangle$.
Appendix B Equations of motion and time evolution scheme
The explicit form of equations of motion (EoMs) for $\Psi_{I}$ and $A_{\mu}$ are
@@ Appendix B Equations of motion and time evolution scheme
For time evolution, we use fourth order Runge-Kutta method and the following scheme: First, we use (32), (34) and (35) to evolve $\Phi$, $A_{x}$ and $A_{y}$ with boundary conditions $\Phi(z=0)=A_{x}(z=0)=A_{y}(z=0)=0$. Then we use (33) to evolve $\partial_{z}A_{t}$ on the boundary. Note that $-\partial_{z}A_{t}(z=0)$ is just the charge density or number density $\rho$ of the dual field theory. Finally, we use (36) to solve $A_{t}$ by evolved $\Phi$, $A_{x}$, $A_{y}$ and boundary conditions $\partial_{z}A_{t}(z=0)=-\rho$ and $A_{t}(z=0)=\mu$. Such scheme keeps $\mu$, the chemical potential, unchanged, so we are in fact working in grand canonical ensemble. In $z$ direction, we use Chebyshev pseudo spectral method. In $x$ and $y$ direction we use Fourier pseudo spectral method and periodic boundary condition.
Appendix C Stationary miscible binary superfluids
@@
$\begin{split}&\Psi_{I}=z\phi_{I}(z)e^{i\Theta_{I}(z,x,y)},\\ &A_{t}=A_{t}(z),\\ &V_{Ix}(z)\equiv\partial_{x}\Theta_{I}-A_{x}\,,\\ &V_{Iy}(z)\equiv\partial_{y}\Theta_{I}-A_{y},\\ &(I=1,2)\,,\\ \end{split}$ (39)
@@
$\begin{split}\partial_{z}(f\partial_{z}\phi_{I})-(V_{Ix}^{2}+V_{Iy}^{2})\phi_{I}+\frac{A_{t}^{2}}{f}\phi_{I}\\ -z\phi_{I}-\frac{\nu}{2}\phi_{J}^{2}\phi_{I}=0,\,(I\neq J)\,,\\ f\partial_{z}^{2}A_{t}-2A_{t}\sum_{I}\phi_{I}^{2}=0\,,\\ \partial_{z}(f\partial_{z}V_{Ix})-2\sum_{I}V_{ix}\,\phi_{I}^{2}=0\,,\\ \partial_{z}(f\partial_{z}V_{Iy})-2\sum_{I}V_{Iy}\,\phi_{I}^{2}=0\,,\end{split}$ (40)
@@
Apparently, there exists two kinds of solution. One is single component superfluid solution $\phi_{I}\neq 0$, $\phi_{J}=0$, and the other is binary superfluid solution, which is the focus of this work. For fixed temperature, binary superfluid solution only exists for $\nu>\nu_{c}$ Zhao et al. (2023). Below this critical value, superfluid would collapse due to strong attraction between the two components. For counterflow case $v_{+}=0$ and coflow case $v_{-}=0$, there is a additional symmetry $\phi_{1}=\phi_{2}=\phi$ between the two components. Utilizing the symmetry, the EoMs (40) can be slightly further simplified. For counterflow binary superfluids $(v_{1})_{y}=-(v_{2})_{y}=v_{y}$, the difference of the two equations for $\phi_{I}$ gives $v_{y}A_{y}\phi=0$, which means $A_{y}=0$ for nonzero $\phi$. Then, the equation of motion for $V_{Iy}$ are eliminated, and the full EoMs become
@@
$\begin{split}\partial_{z}(f\partial_{z}\phi)-V_{y}^{2}\phi+\frac{A_{t}^{2}}{f}\phi-z\phi-\frac{\nu}{2}\phi^{3}=0\,,\\ f\partial_{z}^{2}A_{t}-4A_{t}\phi^{2}=0\,,\\ \partial_{z}(f\partial_{z}V_{y})-4V_{y}\,\phi_{I}^{2}=0\,,\end{split}$ (43)
@@
C.1 Free energy density and susceptibilities
Given the stationary solutions, we can proceed to calculate the free energy density and susceptibilities. According to the holographic dictionary, the free energy is simply identified with on-shell bulk action with Euclidean signature times temperature. In the probe limit with sources of condensates set to 0, we can neglect the gravity part of the action, the Euclidean action is finite and no counter term is needed. Therefore the free energy density can easily be calculated as
[…]
where $\mathcal{L}$ is given by (29). This is essentially the free energy density difference with respect to pure AdS planar black hole without matter fields. To get $\chi_{AB}$, which is given by
$\chi_{AB}=\frac{\delta^{2}f}{\delta s_{A}\delta s_{B}}\,,$ (45)
we need to calculate the free energy density for different $s_{A}$ and $s_{B}$. In this work, $f$ depends on three parameters $s=(\mu,v^{s}_{+},v^{s}_{-})$. Once we get $f(\mu,v^{s}_{+},v^{s}_{-})$, we can use finite difference method to extract the value of $\chi_{AB}$ numerically. For instance, if we want to calculate $\chi_{++}$, we need to calculate free energy density for different $v^{s}_{+}$ while keeping $(\mu,v^{s}_{-})$ fixed to $(\mu_{0},v^{s}_{-0})$, and then calculate $\partial_{v^{s}_{+}}^{2}f(\mu_{0},v^{s}_{+},v^{s}_{-0})$ numerically.
Appendix D Linearized equations of motion
@@
There are some general features of this generalized eigenvalue problem. Two types of symmetries lie in the EoMs. The first one links EoMs for $k$ and $-k$. If $u_{k}$ and $\omega_{k}$ is a solution for (54), we would also have
@@
Therefore for counterflow case, $\omega_{k}$ and $-\omega_{k}^{*}$ comes as pairs for given $k$. If $\omega_{k}=-\omega_{k}^{*}$, i.e. $\omega_{k}$ is pure imaginary, we would also have $(u_{1},v_{1})=(v_{2},u_{2})^{*}e^{i\alpha}$, where $\alpha$ is a global phase. For coflow case, we have $u_{k}=u^{+}_{k}$, which gives $(u_{1},v_{1})=(u_{2},v_{2})e^{i\alpha}=(u_{1},v_{1})e^{2i\alpha}\Rightarrow e^{i\alpha}=\pm 1$. Therefore, there are two sets of solutions $(u_{1},v_{1})=\pm(u_{2},v_{2})$ for coflow case. In the main text, we see that one of the solutions shows Landau instability and the other shows coflow instability.
Appendix E Sound velocity from QNMs and susceptibilities
@@
$(v_{s})_{\pm}=\frac{-\chi_{n+}\pm\sqrt{\chi_{n+}^{2}+\chi_{++}\chi_{nn}}}{\chi_{nn}}\,,$ (58)
with $\chi_{AB}=\delta^{2}f/\delta s_{A}\delta s_{B}$. Sound velocity extracted from these two methods should coincide with each other. As a self consistence check, below in Figure 10, we plot the discrepancy between sound velocities from the two methods $\Delta v_{s}/(v_{s})_{\mathrm{QNM}}=((v_{s})_{\mathrm{QNM}}-(v_{s})_{\mathrm{susceptibility}})/(v_{s})_{\mathrm{QNM}}$ for the larger sound velocity of counterflow and coflow cases separately. We see the results agree well within numerical error.
Figure 10: Discrepancy between sound velocities extracted from QNMs and susceptibilities, $\Delta v_{s}/(v_{s})_{\mathrm{QNM}}=((v_{s})_{\mathrm{QNM}}-(v_{s})_{\mathrm{susceptibility}})/(v_{s})_{\mathrm{QNM}}$, for different superfluid velocities. Left: discrepancy of the larger sound velocity of counterflow case. Right: discrepancy of the larger sound velocity of coflow case. Relavant parameters are $T/T_{c}=0.677$ and $\nu=-0.2$.

### v2

@@ Appendix C Holographic setup
To investigate holographic superfluids with spontaneously broken symmetry, we turn off electromagnetic fields and scalar sources at the boundary, meaning $a_{x}=a_{y}=(\Psi_{I})_{0}=0$. Consequently, the charge conservation equation at the boundary is given by $\partial_{t}b_{t}=b_{i}$, where $b_{t}=n$ represents the charge density and $b_{i}=j_{i}$ is the charge current. The superfluid velocity is determined by $\bm{v}^{s}=\nabla\theta|_{z=0}$, where $\theta$ is the phase (the Goldstone field) of the order parameter $\langle\mathcal{\hat{O}}\rangle$.
Appendix D Equations of motion and time evolution scheme
The explicit form of EoMs for $\Psi_{I}$ and $A_{\mu}$ are
@@ Appendix D Equations of motion and time evolution scheme
For time evolution, we use fourth order Runge-Kutta method and the following scheme: First, we use (52), (54) and (55) to evolve $\Phi$, $A_{x}$ and $A_{y}$ with boundary conditions $\Phi(z=0)=A_{x}(z=0)=A_{y}(z=0)=0$. Then we use (53) to evolve $\partial_{z}A_{t}$ on the boundary. Note that $\partial_{z}A_{t}(z=0)$ is just the charge density or number density $\rho$ of the dual field theory. Finally, we use (56) to solve $A_{t}$ by evolved $\Phi$, $A_{x}$, $A_{y}$ and boundary conditions $\partial_{z}A_{t}(z=0)=-\rho$ and $A_{t}(z=0)=\mu$. Such scheme keeps $\mu$, the chemical potential, unchanged, so we are in fact working in grand canonical ensemble. In $z$ direction, we use Chebyshev pseudo spectral method. In $x$ and $y$ direction we use Fourier pseudo spectral method and periodic boundary condition.
Appendix E Stationary miscible binary superfluids
@@
[…]
@@
[…]
@@
Apparently, there exists two kinds of solution. One is single component superfluid solution $\phi_{I}\neq 0$, $\phi_{J}=0$, and the other is binary superfluid solution, which is the focus of this work. For fixed temperature, binary superfluid solution only exists for $\nu>\nu_{c}$ Zhao et al. (2023). Below this critical value, superfluid would collapse due to strong attraction between the two components. For the counterflow case $v_{+}=0$ and the coflow case $v_{-}=0$, there is a additional symmetry $\phi_{1}=\phi_{2}=\phi$ between the two components. Utilizing the symmetry, the EoMs (60) can be slightly further simplified. For counterflow binary superfluids $(v_{1})_{y}=-(v_{2})_{y}=v_{y}$, the difference of the two equations for $\phi_{I}$ gives $v_{y}A_{y}\phi=0$, which means $A_{y}=0$ for nonzero $\phi$. Then, the equation of motion for $V_{Iy}$ are eliminated, and the full EoMs become
@@
[…]
@@
E.1 Free energy density and susceptibilities
Given the stationary solutions, we can proceed to calculate the free energy density and susceptibilities. According to the holographic dictionary, the free energy is simply identified with on-shell bulk action with Euclidean signature times temperature. In the probe limit with sources of condensates set to 0, we can neglect the gravity part of the action, the Euclidean action is finite and no counter term is needed. Therefore the free energy density can easily be calculated as [^20] Due to our boundary conditions, it is $F$ where $v^{s}_{\pm}$ are held fixed which we calculate, and not $\bar{F}$ where $\bf\bar{h}_{\pm}$} are held fixed.
$f\equiv\frac{F}{Vol}=-\int dz\sqrt{-g}\mathcal{L}_{\mathrm{on-shell}}\,.$ (64)
where $Vol$ is the volume of the spatial subspace on the boundary, $\mathcal{L}$ is given by (49) and $T=3/(4\pi z_{h})$. This is essentially the free energy density difference with respect to pure AdS planar black hole without matter fields. To get $\chi_{AB}$, we need to calculate the free energy density for different values of the parameters $(\mu,v^{s}_{+},v^{s}_{-})$. Once we obtain $f(\mu,v^{s}_{+},v^{s}_{-})$, we can use the finite difference method to extract the value of $\chi_{AB}$ numerically. For instance, if we want to calculate $\chi_{++}$, we need to calculate the free energy density for different $v^{s}_{+}$ while keeping $(\mu,v^{s}_{-})$ fixed to $(\mu_{0},v^{s}_{-0})$, and then calculate $\chi_{++}=-\partial_{v^{s}_{+}}^{2}f(\mu_{0},v^{s}_{+},v^{s}_{-0})$ numerically.
Appendix F Linearized equations of motion
F.1 Linearized EoMs of Hydrodynamics
After inserting the constitutive relations (11) and (13), the explicit form of EoMs (10) is
$\partial_{t}n+\partial_{i}\sum_{I}\bar{h}_{I}^{i}$ (65)
$=\partial_{i}[(\sigma_{1}\delta^{ij}+\sigma_{2}\frac{v_{J}^{si}v_{J}^{sj}}{|v_{J}^{s}|^{2}})\partial_{j}\mu+\beta v_{J}^{si}\partial_{j}\bar{h}_{J}^{j}]\,$,
$\partial_{t}v_{I}^{si}+\partial_{i}\mu$
$=\partial_{i}[\beta v_{I}^{sj}\partial_{j}\mu+(\delta_{1}\delta_{IJ}+\delta_{2}v_{I}^{si}v_{Ji}^{s})\partial_{j}\bar{h}_{J}^{j})]\,.$
By perturbing the expectation values $O_{A}=O_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta O_{A}$ and their sources $s_{A}=s_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta s_{A}$, where $O_{A}=(n,\bm{v^{s}_{I}})$ and $s_{A}=(\mu,\bm{\bar{h}_{I}})$, we can obtain the linearized EoMs
$i\omega\delta n+ik\sum_{I}\delta\bar{h}_{I}$ (66)
$=-(\sigma_{1}+2\sigma_{2})k^{2}\delta\mu-\beta v_{J}^{s}k^{2}\delta\bar{h}_{J}\,$,
$i\omega\delta v_{I}^{s}+ik\delta\mu$
$=-\beta v_{I}^{sj}k^{2}\delta\mu-(\delta_{1}\delta_{IJ}+\delta_{2}v_{I}^{s}v_{J}^{s})k^{2}\delta\bar{h}_{J}\,.$
Changing variables to (16) we get
$\left(\begin{array}[]{c}\delta n\\ \delta v^{s}_{+}\\ \delta v^{s}_{-}\end{array}\right)=M\cdot\left(\begin{array}[]{c}\delta\mu\\ \delta\bar{h}_{+}\\ \delta\bar{h}_{-}\end{array}\right)\,$, (67)
with matrix $M$ given by (17) in the main text.
F.2 Linearized EoMs of Holographic binary superfluids
@@
There are some general features of this generalized eigenvalue problem. Two types of symmetries lie in the EoMs. The first one links EoMs for $k$ and k-k. If $u_{k}$ and $\omega_{k}$ is a solution for (76), we would also have
@@
Therefore for the counterflow case, $\omega_{k}$ and $\omega_{k}^{*}$ come as pairs for given $k$. If $\omega_{k}=-\omega_{k}^{*}$, i.e. $\omega_{k}$ purely imaginary, we would also have $(u_{1},v_{1})=(v_{2},u_{2})^{*}e^{i\alpha}$, where $\alpha$ is a global phase. For the coflow case, we have $u_{k}=u^{+}_{k}$, which gives $(u_{1},v_{1})=(u_{2},v_{2})e^{i\alpha}=(u_{1},v_{1})e^{2i\alpha}\Rightarrow e^{i\alpha}=\pm 1$. Therefore, there are two sets of solutions $(u_{1},v_{1})=\pm(u_{2},v_{2})$ for the coflow case. In the main text, we see that one of the solutions shows the Landau instability and the other shows the coflow instability.
Appendix G Sound velocity from QNMs and susceptibilities
@@
$(v_{s})_{\pm}=\frac{-\chi_{n+}\pm\sqrt{\chi_{n+}^{2}+\chi_{++}\chi_{nn}}}{\chi_{nn}}\,.$ (80)
The sound velocity extracted from these two methods should be in agreement. As a self-consistency check, below in Figure 10, we plot the discrepancy between sound velocities from the two methods $\Delta v_{s}/(v_{s})_{\mathrm{QNM}}=((v_{s})_{\mathrm{QNM}}-(v_{s})_{\mathrm{susceptibility}})/(v_{s})_{\mathrm{QNM}}$ for the larger sound velocity of the counterflow and coflow cases separately. We find the results agree well within numerical error.
Figure 10: Discrepancy between sound velocities extracted from QNMs and susceptibilities, $\Delta v_{s}/(v_{s})_{\mathrm{QNM}}=((v_{s})_{\mathrm{QNM}}-(v_{s})_{\mathrm{susceptibility}})/(v_{s})_{\mathrm{QNM}}$, for different superfluid velocities. Left: discrepancy of the larger sound velocity of the counterflow case. Right: discrepancy of the larger sound velocity of the coflow case. Relevant parameters are $T/T_{c}=0.677$ and $\nu=-0.2$.